### v2

# Shallow IQP circuits and graph generation

Abstract
We introduce shallow instantaneous quantum polynomial-time (IQP) circuits as generative graph models, using an edge-qubit encoding to map graphs onto quantum states. Focusing on bipartite and Erdős–Rényi distributions, we study their expressivity and robustness through simulations and large-scale experiments. Noiseless simulations of $28$ qubits (88-node graphs) reveal that shallow IQP models can learn key structural features, such as the edge density and bipartite partitioning. On IBM’s Aachen QPU, we scale experiments from $28$ to $153$ qubits (881818 nodes) in order to characterize performance on real quantum hardware. Local statistics—such as the degree distributions—remain accurate across scales with total variation distances ranging from $0.04$ to $0.20$, while global properties like strict bipartiteness degrade at the largest system sizes ($91$ and $153$ qubits). Notably, spectral bipartivity—a relaxation of strict bipartiteness—remains comparatively robust at higher qubit counts. These results establish practical baselines for the performance of shallow IQP circuits on current quantum hardware and demonstrate that, even without error mitigation, such circuits can learn and reproduce meaningful structural patterns in graph data, guiding future developments in quantum generative modeling for the NISQ era and beyond.
I Introduction
Figure 1: Overview of the proposed workflow. Top: Model training is performed entirely on classical hardware via a classical estimation of the MMD loss. Bottom: Sampling is carried out on IBM’s Aachen quantum processor.
Graphs represent relationships between entities and provide a unified framework to describe complex systems via their components and interactions [newmanNetworks2018]. Their versatility has driven the development of graph generation methods for applications ranging from drug discovery to scheduling [drobyshevskiyRandomGraphModeling2020, bonifatiGraphGeneratorsState2021, bonginiMolecularGenerativeGraph2021, yangMoleculeGenerationDrug2024, cordeiroRandomGraphGeneration2010].
Machine learning (ML) methods have emerged as powerful tools for graph generation, mirroring their success in other generative modeling tasks [guoSystematicSurveyDeep2022]. These approaches typically fall into two groups [guoSystematicSurveyDeep2022]: sequential methods, which build graphs autoregressively by adding nodes, edges, or motifs [groverGraphiteIterativeGenerative2019, tranDeepNCDeepGenerative2020, bacciuEdgebasedSequentialGraph2020], and one-shot methods, which generate entire graphs in a single pass by sampling from a learned distribution [simonovskyGraphVAEGenerationSmall2018, flam-shepherdGraphDeconvolutionalGeneration2020, niuPermutationInvariantGraph2020]. Sequential methods provide finer control over generation but incur higher sampling costs; one-shot methods trade control for speed.
Nevertheless, learning a graph distribution remains challenging because node interactions span both local and global dependencies. As the number of nodes increases, the combinatorial complexity of pairwise and higher-order relationships grows rapidly: adding a single node to an $M$-node graph introduces $M$ new potential edges, so that the number of possible configurations increases from $2^{\binom{M}{2}}$ to $2^{\binom{M}{2}}\cdot 2^{M}$. This exponential scaling implies that even modest increases in graph size can dramatically expand the hypothesis space. Although practical problems typically restrict attention to structured graph families, models must still possess sufficient expressivity to navigate this high-dimensional space.
Quantum generative modeling emerges as an alternative framework to overcome the limitations of its classical counterpart by leveraging quantum phenomena [zoufalGenerativeQuantumMachine2021]. By encoding information in quantum states, quantum generative models (QGMs) operate in a Hilbert space whose dimension grows exponentially with the number of qubits [nielsenQuantumComputationQuantum2010], allowing them to represent probability distributions that classical systems cannot feasibly store or manipulate. Consequently, QGMs can efficiently learn and reproduce distributions that are intractable for classical models [swekeQuantumClassicalLearnability2021, huangGenerativeQuantumAdvantage2025]. Whether such beyond-classical expressivity yields practical benefits for structured domains such as graph generation remains an open question.
A prominent example of a QGM is the quantum circuit Born machine (QCBM), which encodes probability distributions via the measurement statistics of a parameterized quantum circuit [coyleBornSupremacyQuantum2020]. QCBMs can be implemented using an instantaneous quantum polytime (IQP) ansatz [shepherdInstantaneousQuantumComputation2009], an architecture that has emerged as a promising candidate for quantum generative modeling [recio-armengolTrainClassicalDeploy2025, kurkinUniversalityKerneladaptiveTraining2025, zouConQuERModularArchitectures2025]. These circuits are conjectured to be classically intractable [marshallImprovedSeparationQuantum2024], yet certain statistical properties of their output distributions can be efficiently estimated on classical hardware. This feature enables a hybrid learning framework where training relies on classical estimated statistics, while sampling requires quantum hardware [recio-armengolTrainClassicalDeploy2025].
In this work, we investigate shallow IQP QCBMs for generating random bipartite and Erdős–Rényi graph distributions. These families offer an ideal testbed: their well-defined combinatorial structure and varied features probe different facets of a model’s expressive power. We propose a single-shot graph-generation workflow that trains on classical hardware and runs on current quantum devices (Fig. 1). Classical simulations up to $28$ qubits validate the approach and assess its noise resilience. These show that our models can reproduce the key defining properties of each graph family, with errors that depend on feature complexity and locality.
We study the scaling by deploying $28$-, $45$-, $91$-, and $153$-qubit models on IBM’s Aachen QPU. For global features like bipartiteness, the models outperform the null-hypothesis baseline up to $45$ qubits, demonstrating their ability to reproduce the bipartite structure on real hardware; performance then declines at larger scales owing to increased noise and complexity. A relaxed measure of bipartiteness, spectral bipartivity, continues to outperform the baseline up to $91$ qubits. Notably, even at $153$ qubits, local features such as the degree distribution remain accurately captured. Finally, maximum mean discrepancy tests confirm that the learned distributions remain faithful to the targets.
The paper is organized as follows. Section II defines the target graph distributions, their representations, and our features of interest. Section III describes the proposed shallow IQP model and its optimization. Section IV outlines the experimental setup, including the datasets, training, hyperparameter optimization, and sampling. Section V reports the results. Finally, section VI concludes and discusses future directions.
II Graphs and Random Graph Distributions
Figure 2: Top: Adjacency matrix representation of a 77-node undirected graph. Bottom: Corresponding bit string encoding constructed from the upper-triangular part of the adjacency matrix.
A graph distribution defines a probability distribution over the space of graphs satisfying a given criterion. In this work, we restrict attention to undirected, unweighted graphs to isolate structural dependencies while maintaining a compact binary representation. Such graphs connect nodes by undirected edges: if node $i$ is linked to node $j$, then node $j$ is equally linked to node $i$. Consequently, the adjacency matrix $\mathbf{A}$ is symmetric, $A_{ij}=A_{ji}$, and self-loops are excluded, i.e. $A_{ii}=0\ \forall\ i$.
Because of this symmetry, specifying one triangular portion of $\mathbf{A}$ suffices. An undirected graph with $M$ nodes is therefore fully characterized by
$N=\frac{M(M-1)}{2},$ (1)
binary variables that encode the presence or absence of edges between distinct nodes (see Fig. 2). This encoding defines the effective dimensionality of the generative problem.
We focus on two families of random graph distributions. The first is the Erdős–Rényi (ER) model [erdosRandomGraphs2022], which defines graphs in which each possible edge is included independently with probability $\rho\in[0,1]$, known as the edge density. The second family comprises bipartite (BP) graphs, whose vertices can be partitioned into two disjoint sets such that edges exist only between and never within—equivalently, 2-colorable graphs. BP distributions impose structural constraints that contrast with the edge-independent probabilistic structure of ER graphs, making them a useful testbed.
These two classes are not mutually exclusive: an ER sample may be bipartite, particularly at low edge densities, while any BP graph can be viewed as a constrained instance of the ER model. Throughout this work, BP distributions refer to ensembles of random bipartite graphs, whereas ER distributions denote ensembles of random graphs that may or may not exhibit bipartite structure.
II.1 Bit string graph representation
To map the circuit output to a graph, we flatten the upper triangular part of the adjacency matrix row by row into a binary string $\mathbf{z}=(z_{1},\dots,z_{N})$, where N=M(M1)/2N=M(M-1)/2 and $M$ is the number of nodes, as illustrated in Fig. 2. In this representation, each qubit in the quantum circuit corresponds to a potential edge, and a measurement in the computational basis yields a binary string that uniquely defines a specific graph.
This encoding scales quadratically with the number of nodes. While more compact latent encodings or reparameterizations could reduce the required qubit count, we adopt this direct mapping as an interpretable baseline. It ensures an unambiguous correspondence between quantum measurement outcomes and graph structures, avoiding potential confounding factors introduced by latent representations.
II.2 Features of interest
Graph features represent different degrees of node interactions: some depend only on local edge probabilities, while others require knowledge of global relationships.
We formalize this notion as follows. Let $P(\mathbf{z})$ denote the target distribution over adjacency vectors, $\mathbf{z}=(z_{1},\ldots,z_{N})\in\{0,1\}^{N}$, and let $f:\{0,1\}^{N}\to\mathbb{R}$ be a feature of interest.
Definition 1 ($k$-bodied feature).
A feature $f$ is $k$-bodied if its expected value under $P$, $\mathbb{E}_{P}[f(\mathbf{z})]$, depends only on correlations among at most $k$ edges.
Formally, for any subset of edge indices $S\subseteq\{1,\ldots,N\}$ with $|S|\leq k$, let $P_{S}(\mathbf{z}_{S})$ denote the marginal distribution of the subvector $\mathbf{z}_{S}=(z_{i}:i\in S)$. Then $f$ is $k$-bodied if the collection of all such marginals,
$\big\{P_{S}(\mathbf{z}_{S}):S\subseteq\{1,\ldots,N\},\,|S|\leq k\big\},$ (2)
is sufficient to uniquely determine $\mathbb{E}_{P}[f(\mathbf{z})]$.
Definition 2 (Global feature).
A feature $f$ is said to be global if it is not $k$-bodied for any fixed $k<N$.
These definitions measure the correlations required to reproduce any given property in an idealized setting. We now describe the features used in this work, their theoretical bodyness, and practical considerations.
The simplest feature is density, $\rho$, which measures the fraction of present edges relative to the maximum number of edges,
$\rho(\mathbf{z})=\frac{2}{M(M-1)}\sum_{i=1}^{N}z_{i}.$ (3)
Density provides a coarse summary of the overall connectivity, showing whether a graph is sparse ($\rho\approx 0$) or dense ($\rho\approx 1$), but it does not reveal how edges are arranged among nodes. Since the expected density depends only on single-edge probabilities, it is a 11-bodied feature. Consequently, it should be easy to learn and reproduce.
The next feature of interest is the degree distribution, which captures the number of edges incident on each node and provides a more detailed view of connectivity than density alone. For a node $v_{j}$, the degree is defined as
$\deg(v_{j})=\sum_{i\in\mathcal{E}_{j}}z_{i},$ (4)
where $\mathcal{E}_{j}$ is the set of edges incident on $v_{j}$. The degree distribution of a graph is then the histogram of $\deg(v_{j})$ over all nodes $j\in\{1,\dots,M\}$. For random graphs with independent edges, the degree of a node is the sum of independent Bernoulli variables, yielding a binomial distribution. Therefore, the degree distribution is also 11-bodied. Although it has the same bodyness as the density, it should be harder to reproduce in practice, as it encodes additional information beyond the mean edge probability.
The last feature we consider is the bipartite property, which indicates whether a graph is 22-colorable—i.e., has no odd-length cycles. The degree of correlations required to determine this property increases with graph size; for an $M$-node graph, one must capture all odd-cycle correlations up to length MM—or M1M-1 if $M$ is even. Consequently, bipartiteness is a global feature, as it depends on correlations that span the entire graph.
Among the considered features, bipartiteness is the most difficult to reproduce. It’s a binary property—a graph is either bipartite or not—plus it is highly sensitive to perturbations: even a single bit flip can introduce an odd-length cycle, converting a bipartite graph into a non-bipartite one.
III IQP circuit born machines
A QCBM is a quantum generative model that represents a probability distribution in the measurement statistics of a pure quantum state [liuDifferentiableLearningQuantum2018]. Sampling from this distribution is achieved through projective measurements in the computational basis, producing single-shot samples.
In this work, we employ QCBMs constructed with IQP circuits. These represent a class of quantum computations that are restricted to commuting operations [shepherdInstantaneousQuantumComputation2009, nakataDiagonalQuantumCircuits2014]. Despite this apparent limitation, they have emerged as a powerful tool for generative modeling. Sampling from most instances of IQP circuits is conjectured to be classically hard [bremnerAverageCaseComplexityApproximate2016], as simulating them efficiently would imply a collapse of the polynomial hierarchy to its second level [bremnerClassicalSimulationCommuting2011, marshallImprovedSeparationQuantum2024]. Yet, classical computers can efficiently estimate the expectation values of their commuting operators [nestSimulatingQuantumComputers2010]. This feature enables their classical optimization when the cost function depends solely on those values [recio-armengolIQPoptFastOptimization2025]. Thus, it allows the implementation of a hybrid framework where classical resources are used for training and quantum resources for sampling.
Our objective is to train shallow IQP models to generate graphs according to a ground truth distribution $p$. Specifically, we aim to learn a parameterized distribution $q_{\bm{\theta}}$ such that, for any feature of interest $f(\mathbf{z})$,
$\mathbb{E}_{\mathbf{z}\sim q_{\bm{\theta}}}[f(\mathbf{z})]-\mathbb{E}_{\mathbf{z}\sim p}[f(\mathbf{z})]<\varepsilon,$ (5)
where $\varepsilon$ defines the target accuracy.
III.1 Shallow IQP circuits
The foundation of our models is a NN-qubit shallow IQP circuit, building on the general definition from Nakata et al. [nakataDiagonalQuantumCircuits2014].
Definition 3.
A shallow IQP circuit on $N$ qubits is defined by the following sequential operations,
  1. 1.
    Basis Transformation: An initial layer of Hadamard gates $H^{\otimes N}$ applied to all qubits.
  2. 2.
    Diagonal Evolution: A constant depth parameterized block of gates $D_{Z}(\bm{\theta})$ that is diagonal in the Pauli-Z basis. This block is composed of single- and two-qubit rotations generated by Pauli-Z operators.
  3. 3.
    Inverse Transformation: A final layer of Hadamard gates $H^{\otimes N}$ applied to all qubits.
  4. 4.
    Measurement: A projective measurement in the computational basis.
The circuit diagram form is illustrated in Figure 3.
This sequence of operations implements a unitary $U(\bm{\theta})=H^{\otimes N}D_{Z}(\bm{\theta})H^{\otimes N}$, where $\bm{\theta}=\{\theta_{i}\}_{i=1}^{n}$, that prepares the following quantum state,
$|\Psi(\bm{\theta})\rangle=H^{\otimes N}D_{Z}(\bm{\theta})H^{\otimes N}|0\rangle^{\otimes N},$ (6)
which encodes a probability distribution with respect to measurements in the computational basis.
Crucially, for any observable $O_{z}$ that is a product of Pauli-Z operators, its expectation value,
$\langle\Psi(\bm{\theta})|O_{z}|\Psi(\bm{\theta})\rangle,$ (7)
can be computed efficiently on classical hardware. The precise complexity of this process is detailed in Proposition 2 of Recio-Armengol et al. [recio-armengolIQPoptFastOptimization2025].
Figure 3: Circuit diagram of a parameterized IQP circuit. The unitary $D_{Z}(\bm{\theta})$ contains all diagonal parameterized gates. For constant-depth $D_{Z}(\bm{\theta})$, the circuit realizes a shallow instance of the IQP model.
All parameterized operations are contained within the diagonal gate block. Therefore, the ansatz is fully specified by this block alone. Crucially, it determines both the expressivity and the hardware efficiency of a parameterized circuit. A simple ansatz may be easy to implement but fail to capture high-bodied correlations, whereas a more complex one may be able to represent such correlations but at the cost of increased resource requirements—and an increased noise sensitivity. To balance this trade-off, we employ a shallow ansatz that uses few resources while still being conjectured to be classically intractable.
The classical simulability of IQP circuits depends on the type and connectivity of the gates employed. Circuits composed solely of single-qubit rotations are trivially simulable, as they generate no entanglement. Furthermore, circuits restricted to nearest-neighbor two-qubit gates can also be efficiently simulated [fujiiQuantumCommutingCircuits2017]. Notably, circuits that combine these two layouts fall out of the simulable regime while maintaining a shallow depth [fujiiQuantumCommutingCircuits2017]. This is the chosen design for this work, one of the simplest IQP instances conjectured to be classically intractable.
This shallow ansatz maps efficiently to the target quantum processor (QPU), IBM’s Aachen, since both parameterized gates belong to the device’s native gate set. In particular, RZ gates are implemented as virtual rotations with negligible latency and error [mckayEfficientZGatesQuantum2017]. Although Hadamard gates are not native, the compiler decomposes them into RZ($\pi/2$) and SX gates, introducing only a minimal depth overhead.
Notably, we omit ancilla qubits, even though they are necessary to achieve universality [kurkinNoteUniversalityParameterized2025]. While this choice limits expressive power, it may also enhance trainability—as expressivity and barren plateaus are closely linked [laroccaBarrenPlateausVariational2025, cerezoDoesProvableAbsence2025].
III.2 Optimization via the maximum mean discrepancy
To train an IQP model on classical hardware, one must formulate the optimization task in terms of expectation values—both from the target data and the circuit itself.
The maximum mean discrepancy (MMD) can be formulated in such terms. It is an integral probability metric that compares two probability distributions, $p$ and $q$, based on samples, $\bm{x}$ and $\bm{y}$, drawn from each in a reproducing kernel Hilbert space (RKHS) [JMLR:v13:gretton12a]. The squared MMD is defined as [sutherlandGenerativeModelsModel2021]
$\begin{split}MMD^{2}(p,q_{\theta})&=\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim p}[k(\bm{x},\bm{y})]\\ &\quad-2\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})]\\ &\quad+\mathbb{E}_{\bm{x}\sim q_{\theta},\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})],\end{split}$ (8)
where $k(\bm{x},\bm{y})$ is a positive definite kernel. If $k$ is characteristic, then $MMD^{2}=0$ if and only if $p=q_{\theta}$ [JMLR:v13:gretton12a]. The three terms respectively measure intra-data similarity, model-data cross-similarity, and intra-model similarity.
Central to this approach, we use the Gaussian kernel,
[…]
for vectors $\bm{x},\bm{y}\in\{0,1\}^{n}$. Since each bit contributes either 0 or $1$ to the sum, the squared Euclidean distance $\|\bm{x}-\bm{y}\|^{2}$ coincides with the Hamming distance, i.e., the number of positions at which the entries of the vectors differ. This equivalence allows the MMD to be expressed in terms of Pauli-Z observables [rudolphTrainabilityBarriersOpportunities2024],
$MMD^{2}(\bm{\theta})=\mathbb{E}_{\bm{a}\sim\mathcal{P}{\sigma}(\bm{a})}\left[(\langle Z_{\bm{a}}\rangle_{p}-\langle Z_{\bm{a}}\rangle_{q_{\bm{\theta}}})^{2}\right],$ (10)
where $\bm{a}\in\{0,1\}^{n}$ encodes a ‘mask’ that corresponds to a Pauli-Z observable,
$Z_{\bm{a}}=\prod_{i=1}^{n}Z_{i}^{a_{i}}.$ (11)
The expectation value measures parity correlations among the selected bits,
$\langle Z_{\bm{a}}\rangle_{p}=\mathbb{E}_{\bm{x}\sim p}[(-1)^{\bm{a}\cdot\bm{x}}],$ (12)
where $(-1)^{\bm{a}\cdot\bm{x}}=+1$ for even parity and 1-1 for odd parity.
Crucially, the weighting distribution $\mathcal{P}_{\sigma}$ determines which correlations the loss emphasizes,
$\mathcal{P}_{\sigma}(\bm{a})=(1-p_{\sigma})^{n-|\bm{a}|}(p_{\sigma})^{|\bm{a}|},\quad p_{\sigma}=\frac{1-e^{-1/2\sigma^{2}}}{2}.$ (13)
Each bit appears independently with probability $p_{\sigma}$, so the active bit count $|\bm{a}|$ follows a binomial distribution with mean $np_{\sigma}$ [rudolphTrainabilityBarriersOpportunities2024].
This behavior has a profound impact on training. When $\sigma$ is constant, $\sigma\in O(1)$, the MMD emphasizes high-order correlations, effectively making the loss global [rudolphTrainabilityBarriersOpportunities2024]. Such cost function is known to yield barren plateaus in unstructured circuits, potentially rendering the model untrainable [cerezoCostFunctionDependent2021]. Conversely, if $\sigma$ scales with the number of qubits, $\sigma\in O(n)$, the MMD primarily captures low-order correlations, ensuring non-vanishing gradients and preserving trainability. However, it blinds the loss to the higher-order correlations.
This observation directly connects to the learnability of the target properties. Features such as density are expected to be learnable in a regime free of barren plateaus, whereas bipartiteness may require a constant $\sigma$ and thus lie in an untrainable regime.
Table 1: Summary of the datasets. $\bar{\rho}$ denotes the average density of the samples, BP (%) the fraction of bipartite graphs, $\bar{\beta}$ the average spectral bipartivity, and $N$ the total number of samples.
Nodes Type Dense Medium Sparse
$\bar{\rho}$ BP.% $\bar{\beta}$ N $\bar{\rho}$ BP.% $\bar{\beta}$ N $\bar{\rho}$ BP.% $\bar{\beta}$ N
8 BP $0.3110$ 100 1.0 261 $0.2887$ 100 1.0 271 $0.2256$ 100 1.0 133
ER $0.7618$ 0.0 0.55 200 $0.4420$ 0.5 0.79 200 $0.2207$ 51.5 0.95 200
10 BP $0.3470$ 100 1.0 498 $0.2307$ 100 1.0 500 $0.1771$ 100 1.0 473
ER $0.7884$ 0.0 0.51 500 $0.4380$ 0.2 0.69 500 $0.1919$ 33.8 0.94 500
14 BP $0.3727$ 100 1.0 995 $0.2084$ 100 1.0 999 $0.1042$ 100 1.0 995
ER $0.8145$ 0.0 0.50 1000 $0.4487$ 0.0 0.57 1000 $0.1575$ 27.9 0.93 1000
18 BP $0.3697$ 100 1.0 995 $0.1967$ 100 1.0 998 $0.0747$ 100 1.0 992
ER $0.8255$ 0.0 0.5 1000 $0.4428$ 0.0 0.53 1000 $0.1497$ 17.7 0.89 1000
IV Experimental setup
This section describes the experimental framework used to evaluate the proposed workflow (Fig. 1). We begin by outlining the generation of synthetic graph datasets employed for training and validation. The following subsections detail the model training procedure, hyperparameter optimization (HPO), and sampling strategy. We then introduce the quantitative metrics used to compare the generated and target graph distributions. Together, these components form a reproducible pipeline for assessing the model’s performance.
All the code and data supporting this work are available in the following GitHub repository.
IV.1 Datasets
We generated synthetic datasets for the two graph families introduced in Sec. II: Erdős–Rényi (ER) graphs, where each edge is included independently with probability $\rho$, and bipartite (BP) graphs, where a 22-colorability constraint is imposed. We use the NetworkX library [SciPyProceedings_11] to build them, ensuring that each sample represents a unique, non-isomorphic graph using the procedure described in Sec. II.1.
For each graph family, we make sparse, medium, and dense instances with node counts $M\in\{8,10,14,18\}$, corresponding to vectors of $28$, $45$, $91$, and $153$ bits, respectively. These categories refer to relative densities within a given node count and family; that is, for a fixed $M$ and graph type, sparse instances have lower average density than medium ones, and so on, though absolute values vary across families and sizes. Lastly, the size of each dataset depends on the number of nodes, as smaller graphs admit fewer unique configurations. Table 1 summarizes the properties of each dataset, including graph type, average density, percentage of bipartite graphs, average bipartivity, and number of samples. A detailed description of these metrics is provided in Sec. IV.5.
IV.2 Training
The training objective is to minimize the MMD between the target and generated distributions. We train all models on a single laptop using the IQPopt library [recio-armengolIQPoptFastOptimization2025], an open-source tool for classically optimizing parameterized IQP circuits in JAX. We configure it with the ADAM optimizer [kingmaAdamMethodStochastic2017], the median heuristic for kernel bandwidth selection, and the data-driven initialization method of Recio-Armengol et al. [recio-armengolTrainClassicalDeploy2025], which sets two-qubit gate parameters proportional to the covariance of the corresponding bits in the training data. To refine these initial settings, we introduce scaling hyperparameters for both the kernel bandwidth and the initialization.
IV.3 Hyperparameter optimization
To ensure a consistent basis for comparison across models, we perform HPO using the Optuna library [optuna_2019]. Each configuration is evaluated via repeated $k$-fold cross-validation, with two repetitions and between three and five folds depending on the dataset size.
We optimize three key parameters: the learning rate, which controls the optimizer step size; the bandwidth multiplier, which scales the median-heuristic kernel width; and the initialization multiplier, which adjusts the amplitude of the initial weights. Other settings—such as the number of ansatz layers and the optimizer choice—showed negligible influence on performance in preliminary tests and were therefore kept fixed.
IV.4 Sampling
After training, we generate samples by preparing the $N$-qubit state with the optimized parameters on the IBM Aachen quantum computer and measuring in the computational basis. Each measurement produces an $N$-bit string $\mathbf{z}$, which we map to a graph following the procedure in Sec. II.1.
Importantly, we don’t employ any error mitigation technique or classical post-processing, which allows us to directly assess the raw performance of the quantum hardware.
IV.5 Evaluation metrics
Table 2: Performance validation for the 88-node datasets ($28$-qubit models) conducted via simulations using PennyLane’s lightning.qubit simulator.
Nodes Type Dense Medium Sparse
$\mathbb{E}[\rho]$ (Difference) BP.% (Target %) $\mathbb{E}[\rho]$ (Difference) BP.% (Target %) $\mathbb{E}[\rho]$ (Difference) BP.% (Target %)
8 BP 0.24 (-0.071) 63.09 (100) 0.29 (-0.018) 52.15 (100) 0.225 (-0.0001) 69.34 (100)
ER 0.701 (-0.062) 0.0 (0.0) 0.461 (0.019) 3.1 (0.5) 0.2862 (0.066) 31.5 (51.5)
[…]
Figure 5: Degree distributions obtained from models from $28$ to $153$ qubits trained on Erdős-Rényi datasets and executed on NISQ hardware. Bars indicate empirical node-degree frequencies, while solid lines denote theoretical binomial targets. The total variation distance (TVD) measures the deviation between generated and target distributions.
Figure 6: Scaling of the maximum mean discrepancy (MMD) across all shallow IQP models.
Figure 7: Scaling behavior of bipartite graph generation as a function of empirical density $\rho$ for models trained on bipartite datasets. Curves correspond to different qubit counts, with the dotted line showing the null-hypothesis baseline at each density.
We evaluate model performance by comparing the generated and target graph distributions across the features of interest.
For the density, we measure the deviation in the expected value between the generated and target distributions,
$\Delta\mathbb{E}[\rho]=\mathbb{E}_{\mathbf{z}\sim D_{\bm{\theta}}}[D(\mathbf{z})]-\mathbb{E}_{\mathbf{z}\sim D}[D(\mathbf{z})].$ (14)
This test verifies whether the model reproduces the correct average edge density. Since this reflects the overall edge occupancy, a model that fails here will likely struggle to capture more complex features.
For the degree distribution, we employ a more comprehensive comparison. In an Erdős–Rényi graph, where edges are independent, the degree distribution follows a binomial law. Accordingly, we compute the expected degree distribution from the generated samples and compare it to the corresponding binomial, both qualitatively—via histograms—and quantitatively using the total variation distance (TVD)
$\mathrm{TVD}(p,q_{\bm{\theta}})=\frac{1}{2}\sum_{k}\big|p(k)-q_{\bm{\theta}}(k)\big|.$ (15)
Here $p(k)$ denotes the probability that a randomly selected node has degree $k$, and $q_{\bm{\theta}}(k)$ the corresponding probability under the model with parameters $\bm{\theta}$. This metric captures discrepancies in the full degree distribution, providing a sensitive test of how well local edge correlations are reproduced.
Lastly, we quantify the bipartite structure of the generated distributions using two complementary measures. First, bipartite accuracy, defined as the percentage of bipartite graphs generated, is determined via an exact 22-coloring check. Given that this property is binary and can be very sensitive to noise—a single misplaced edge can introduce an odd-length cycle—we complement it with an expected bipartivity measure based on the spectral bipartivity index $\beta$ introduced by Estrada et al. [estradaSpectralMeasuresBipartivity2005]
$\beta=\frac{\sum_{i}\cosh{\lambda_{i}}}{\sum_{i}e^{\lambda_{i}}},$ (16)
where $\lambda_{i}$ are the eigenvalues of the adjacency matrix $\mathbf{A}$. This index compares the weighted contributions of even and odd length cycles: bipartite graphs yield $\beta=1$, while complete graphs yield $\beta=0$.
We compute these metrics on $512$ samples per trained model. Simulations use PennyLane’s lightning.qubit backend [bergholmPennyLaneAutomaticDifferentiation2022], while hardware experiments are executed on IBM’s $156$-qubit Aachen QPU, with each batch completing in $5.0\pm 0.5$ seconds.
V Results
We organize the analysis in three stages. First, we validate the models at the $28$-qubit scale through noiseless simulations. Next, we assess their noise robustness by comparing the results of identical models from simulations and quantum hardware. Finally, we investigate scaling behavior by deploying models ranging from $28$ to $153$ qubits.
For each model, we perform multiple HPO experiments across different hyperparameter ranges to ensure thorough exploration. From these, we select the best-performing model for each dataset according to the following criteria: for BP datasets, the model that generates the highest proportion of bipartite graphs; and for ER datasets, the model that minimizes the expected density error.
V.1 Model validation
We validate the models using noiseless $28$-qubit simulations to assess their ability to reproduce the expected density and bipartite structure. Table 2 reports the simulated results for the best-performing model on each dataset.
Across ER categories, the models reproduce average edge densities with small deviations: from $0.019$ to $0.066$. Notably, the ER–Sparse model produced $31.5\%$ bipartite graphs; since the ground-truth contained $51.5\%$ bipartites, this corresponds to a $\approx 61.1\%$ accuracy, which is comparable to the performance of BP models.
For the BP families, bipartite accuracy depends on density. To separate learned structure from random occurrence, we use the Erdős–Rényi model as a null-hypothesis baseline: for each dataset, we generate $10^{6}$ random graphs matching the target density and measure the fraction that are bipartite. Performance close to the baseline indicates random occurrence, while performance above the baseline indicates that the model captures bipartite structure beyond random generation. For all three densities, the models exceed the null-hypothesis baseline (Table 3).
Table 3: Comparison of bipartite accuracies (%) from BP models with those from an Erdős–Rényi graph model of equal density. BP accuracies are computed from $512$ generated samples, while the baseline is obtained from $10^{6}$ random realizations.
Density Baseline Generated Improvement
Dense 25.15 63.09 +37.94
Medium 32.16 52.15 +19.99
Sparse 55.82 69.34 +13.52
Overall, the BP–Sparse model attains the best combined performance, with bipartite accuracy $69.34\%$ and an empirical density error of $10^{-4}$ (Table 2). The BP-Medium and BP-Dense models reached $52.15\%$ and $63.09\%$, respectively. The comparatively strong performance of the dense model reflects its underestimation of density, which inadvertently increases the chance of generating bipartite graphs.
These results show that, in noiseless simulation, shallow IQP circuits can learn both the expected density and bipartite structure beyond a null-hypothesis baseline.
V.2 Model noise resilience
We evaluate the noise resilience of our models by comparing their performance on IBM’s Aachen QPU with ideal, noiseless simulations for the $M=8$ case. Consistent outcomes across both backends indicate robustness to hardware noise, while large deviations reveal sensitivity. Through the following sections, we interchangeably use the terms NISQ and quantum hardware to refer to the QPU.
Figure 4(a) shows that the densities measured on NISQ hardware closely match the simulated values, yielding comparable outcomes across all models. Figure 5(a) further shows that the degree distributions measured on the quantum hardware remain in close agreement with the theoretical binomial at the $28$-qubit scale. Together, these findings indicate that shallow IQP models can learn and reproduce single-bodied features on both simulators and quantum hardware.
Bipartite accuracy shows substantially larger degradation on quantum hardware (Fig. 4(b)). The simulated models generated roughly twice as many bipartite graphs as NISQ executions. In contrast, the expected spectral bipartivity (Fig. 4(c)) remains similar across both backends, indicating that relaxations of binary features resist noise more effectively. The MMD comparison (Fig. 4(d)) follows a similar trend: simulations consistently outperform NISQ executions, though the absolute MMD differences remain small, below $0.05$.
Together, these observations show that the inherent noise of NISQ devices degrades the performance in proportion to each feature’s bodyness and noise sensitivity: local, single-body statistics stay robust; relaxations of global properties degrade moderately; and binary global features deteriorate the most. Crucially, at the 88-node scale, noise reduces precision but preserves the learned structure.
V.3 Model scaling
Building on the validation and noise-resilience experiments, we study how shallow IQP models scale towards the limits of current quantum hardware. While the inherent imperfections of NISQ hardware limit the conclusions regarding their expressive power and potential, these results provide a reference point for their performance on current quantum devices.
We use the MMD as a global accuracy metric to track how the learned distributions diverge from the targets as the system grows (Fig. 6). Within the ER family, the ER–Medium model achieves the best performance across all sizes, showing the lowest MMD values. ER–Sparse and ER–Dense follow, maintaining MMD levels similar to BP–Sparse and BP–Medium. In contrast, the BP–Dense model performs worst at nearly all scales and degrades sharply at the $153$-qubit level.
For the ER models, we analyze local connectivity through the degree distribution and its TVD from the theoretical binomial (Fig. 5). The $28$- and $45$-qubit models reproduce the target distribution accurately, reaching TVDs below $0.055$. Larger systems yield less accurate results. At $91$ qubits, the sparse model underestimates the mean degree, generating overly sparse samples, while the medium and dense models align more closely with the target, giving an average TVD of $0.2038$. At $153$ qubits, the medium and dense models maintain similar performance, whereas the sparse model improves, reducing the average TVD to $0.1012$.
For the BP models, we examine structural properties through their generation accuracy and spectral bipartivity. Accuracy declines as connectivity and qubit count increase, dropping at higher densities and larger scales (Fig. 7), which reflects the growing complexity of this feature. At the $28$- and $45$-qubit scales, the models consistently exceed the baseline accuracy across all densities, whereas at $91$ and $153$ qubits, they generally fall below it. Spectral bipartivity follows the same trend for the smaller systems but performs slightly better overall, with the $91$-qubit models surpassing the baseline more consistently.
These results show that shallow IQP circuits can accurately reproduce low-bodied features even at high qubit counts but struggle with properties that rely on higher-order correlations.
VI Conclusion
In this work, we studied shallow IQP circuits through the task of graph generation with experiments ranging from $28$-qubit simulations to $153$-qubit quantum hardware runs. Our results show that these models can effectively capture low-bodied features at scale, achieving an average TVD of $0.101$ on the generated degree distributions at $153$ qubits. In contrast, higher-bodied features proved to be more complex; however, up to the $45$-qubit scale, the models still surpassed the baseline for bipartite accuracy—the most challenging property we examined.
Furthermore, we employed one of the simplest shallow ansätze conjectured to be classically intractable and applied no error mitigation or post-processing techniques. Such methods would likely enhance performance, as our simulations indicate that reducing noise improves performance, particularly for the most noise-sensitive features. Therefore, these results establish a raw performance baseline for generative graph modeling on current quantum hardware.
Our findings indicate that shallow IQP circuits are most effective for modeling distributions dominated by low-bodied features, in particular at higher qubit counts. For global features, smaller models, despite their limitations, may be better suited as hardware noise will be less noticeable.
IQP-based generative modeling remains a young approach with significant potential. Its classical trainability enables building models that were previously not possible. However, it introduces several challenges: optimizing requires expressing the cost function in terms of expectation values; identifying suitable alternative loss functions remains an open question; and it is not known whether they suffer from barren plateaus. Future work should address these challenges. On the practical side, the impact of error mitigation and post-processing techniques should be explored, as better performance is expected. Furthermore, alternative training schemes, such as adversarial learning, could offer additional advantages and provide deeper insights into how IQP circuits function as generative models.
Acknowledgments
OB and MA are fellows of Eurecat’s “Vicente López” PhD grant program. This piece of research was carried out with the partial support of the following granted projects: SGR Grant 2021 SGR 01559 and RETECH EMT/43/2025 QML-CV from the Catalan Government, GRAIL PID2021-126808OB-I00 and from FEDER/UE, SUKIDI PID2024-157778OB-I00 grants from the Spanish Ministry of Science and Innovation, with the support of Cátedra UAB-Cruïlla grant TSI-100929-2023-2 from the Ministry of Economic Affairs and Digital Transformation of the Spanish Government.